# Automated Black-box Prompt Engineering for Personalized Text-to-Image Generation

### v1

@@ Abstract
Prompt engineering is effective for controlling the output of text-to-image (T2I) generative models, but it is also laborious due to the need for manually crafted prompts. This challenge has spurred the development of algorithms for automated prompt generation. However, these methods often struggle with transferability across T2I models, require white-box access to the underlying model, and produce non-intuitive prompts. In this work, we introduce PRISM, an algorithm that automatically identifies human-interpretable and transferable prompts that can effectively generate desired concepts given only black-box access to T2I models. Inspired by large language model (LLM) jailbreaking, PRISM leverages the in-context learning ability of LLMs to iteratively refine the candidate prompts distribution for given reference images. Our experiments demonstrate the versatility and effectiveness of PRISM in generating accurate prompts for objects, styles, and images across multiple T2I models, including Stable Diffusion, DALL-E, and Midjourney.
Keywords:
Text-to-Image Generation Prompt Engineering Personalized Text-to-Image Generation
@@ 1 Introduction
Figure 1: Given a set of reference images, our method, PRISM, is capable of creating human-interpretable and accurate prompts for the desired concept that are also transferable to both open-sourced and closed-sourced text-to-image models.
@@
An important goal of generative modeling is to design algorithms capable of steering generative models to produce desired output images. Early attempts, which often centered on particular architectures or tasks, were largely characterized by manually-curated data collection, fine-tuning, or retraining from scratch [33, 19, 10, 47]. These requirements are often costly, and the resulting solutions usually do not transfer well between models. Thus despite the promise of these methods, efficient and generalized algorithms for controllable generation remain sought after.
@@
Today, perhaps the most popular approach for controllable generation is to guide the generation process with a piece of textual information, or prompt, that describes the properties of the desired output using text-to-image (T2I) generative models [24, 42]. Through text, T2I models allow users to quickly and easily describe a wide variety of concepts, and model designers can more efficiently explore the behavior of their model through a myriad of strategies [3, 39]. The predominant method for obtaining such input text is to manually design candidate prompts in an iterative, trial-and-error fashion, a process known as prompt engineering, based on what the user (prompt engineer) believes will lead to a desirable output. Unfortunately, these practices are often sensitive to different phrasings [36], require expert domain knowledge, and are notably inefficient as they necessitate a human in the loop.
@@
Motivated by the drawbacks of manual prompt engineering, a recent line of work known as personalized or subject-driven T2I generation has sought to automate the controllable generation pipeline. Given a collection of reference images that capture specific concepts, such as artistic style or shared objects, personalized T2I algorithms are designed to produce images that reflect those concepts illustrated in the reference images. While personalized T2I methods often involve fine-tuning or retraining the underlying T2I model [26, 4, 29], several approaches focus specifically on automating prompt engineering to generate effective prompts. Unfortunately, existing algorithms in this spirit tend to require pre-collected, architecture-specific keywords [^1] https://github.com/pharmapsychotic/clip-interrogator or white-box, embedding-based optimization [6, 16], leading to non-interpretable prompts [39] and precluding the possibility of directly generating prompts for closed-source T2I models (e.g., Midjourney or DALL-E).
@@
In order to address these shortcomings, we propose Prompt Refinement and Iterative Sampling Mechanism (PRISM), a new automated prompt engineering algorithm for personalized T2I generation. A key observation is that prompt engineers repeat the process of updating their “belief” of what makes an effective prompt based on the difference between their desired results and the generated images from previous iterations. Inspired by jailbreaking attacks on large language models (LLMs) [3], we design an algorithm that operates with only limited human input, is capable of generating human interpretable and editable prompts, makes minimal assumptions about the underlying T2I generative model, and generalizes across different T2I models, including popular black-box models such as DALL-E and Midjourney.
@@
Given a set of reference images, our method first generates an initial prompt and its corresponding image using a multimodal LLM and a T2I generative model. We then obtain a score indicating the visual similarity of the generated image and the reference image with respect to the targeting concept via another multimodal LLM. Leveraging LLMs’ in-context learning abilities [30, 46, 40], we instruct the LLM to update the candidate prompt distribution based on the previously generated prompt, images, and the evaluation scores. This processing is then repeated for a predetermined number of iterations. In the end, PRISM outputs the best-performing prompt by re-evaluating the top prompts generated from this process.
@@
Experimentally, our results indicate that PRISM consistently outperforms existing methods, including Textual Inversion [6], PEZ [39], BLIP2 [12] and CLIP-Interrogator1, with respect to human-interpretability while maintaining high visual accuracy. Our method also shows significantly better generalizability and transferability as we achieve the best performance in almost all metrics when experimenting with closed-source models in comparison to baselines. Finally, we also show that because of the interpretability provided by our method, the prompts produced by PRISM are also easily editable, enabling a wide range of creativity possibilities in real life.
@@ 2 Related Works
2.0.1 Controllable T2I generation.
Several methods tackle conditional image generation in a training-free manner by using pretrained diffusion models as priors for the data distribution [32, 18, 5, 31, 9], and analogous approaches exist for T2I diffusion models (e.g., StableDiffusion) [43, 25, 8]. In general, these methods assume that the controllability objective can be formulated as differentiable loss functions, although they require access to model parameters and involve complex hyperparameter tuning. Another class of approaches such as ControlNet [44], IP-Adapter [41], Dreambooth [26], SuTI [4] and InstantBooth [29] also improve the controllability of pretrained T2I models, but they require expensive fine-tuning or re-training of the underlying model. Prompt tuning methods such as Textual Inversion [6], PEZ [39], and PH2P [39, 16] are in the same spirit as this paper, as they do not require fine-tuning or optimizing the underlying model and generate images that inherit the properties of a given reference image. However, unlike PRISM, each of these methods requires access to the underlying model parameters and produces non-interpretable prompts.
2.0.2 Prompt engineering.
Manual prompt engineering is one of the most popular approaches to eliciting desired behaviors from large pre-trained models because it uses little or no data and does not require fine-tuning [22, 2]. However, major drawbacks of manual prompt engineering include its laborious nature, its reliance on domain expertise, and the fact that its performance can be highly sensitive to how the prompts are phrased [15, 36]. To address this issue, several methods have been proposed to construct the prompts in an automated manner [30, 7, 45, 46, 17, 40]. In particular, the field of LLM jailbreaking is concerned with automatically designing prompts that can elicit specific content (which is often objectionable or illicit) from a targeted LLM [48, 37, 23, 14]. A particularly relevant work is [3], which uses an auxiliary LLM to iteratively construct jailbreak prompts that elicit harmful behaviors from a targeted LLM. Our method builds on this idea to generate prompts that will result in images that satisfy the desired criteria.
@@ 3 Method
Figure 2: An illustration of PRISM. The label “System” indicates the system prompts setups for the multimodal LLMs.
@@ 3.1 Problem Statement
First, let $x\in{\mathcal{X}}$ denote an image, and $y\in{\mathcal{Y}}$ denote a textual prompt. Given a collection of reference images $\{x_{i}\}_{i=1}^{M}$, a prompt engineer $\mathbf{F}:{\mathcal{X}}\rightarrow\Delta({\mathcal{Y}})$ samples a candidate prompt $y$ corresponding to each reference image $x$, i.e., $y\sim p_{\theta_{{\mathbf{F}}}}(y\mid x)$. A T2I generative model $\mathbf{G}:{\mathcal{Y}}\rightarrow\Delta({\mathcal{X}})$ then uses this candidate prompt to generate a new image, $x\sim p_{\theta_{\mathbf{G}}}(x\mid y)$, and a judge model $\mathbf{D}:{\mathcal{X}}\times{\mathcal{X}}\rightarrow[0,1]$ then scores the visual similarity between the images based on some criteria. Our goal is then to find the best prompt:
$y^{\star}\left(\{x_{i}\}_{i=1}^{M}\right)=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}\sum_{i=1}^{M}\texttt{Score}(x_{i},y),$ (1)
where $\texttt{Score}(x_{\text{target}},y)=\mathbb{E}_{x\sim p_{\theta_{\mathbf{G}}}(x\mid y)}\left[{\mathbf{D}}(x,x_{\text{target}})\right]$.
@@ 3.2 Algorithm
8:         $\mathbf{D}$ calculates an in-iteration score $\texttt{Score}^{\prime}(x_{k,n},y_{k,n})=\mathbf{D}(x_{k,n},\hat{x}_{k,n})$
9:         Update $p_{\theta_{\mathbf{F}}}$ based on $x_{k,n},\hat{x}_{k,n},y_{k,n},\texttt{Score}^{\prime}(x_{k,n},y_{k,n})$ and the chat history of stream $n$
@@
13:Re-evaluate this subset with total score $\sum_{i=1}^{M}\texttt{Score}(x_{i},y_{c})$
14:Return the prompt with the best total score. In case of a tie, return the prompt with the highest log-likelihood.
Our method, Prompt Refinement and Iterative Sampling Mechanism  (PRISM), is an iterative process that repeats a prompt-refinement subroutine for $K$ iterations in $N$ parallel streams, where $N\times K$ is a predetermined compute budget. At iteration $k$, the $n$-th stream of PRISM  randomly selects a reference image $x_{k,n}$ from $\{x_{i}\}_{i=1}^{M}$ and uses ${\mathbf{F}}$ to sample a candidate prompt $y_{k,n}$ from $p_{\theta_{\mathbf{F}}}(y\mid x_{k,n})$. Then it queries ${\mathbf{G}}$ to generate a single $\hat{x}_{k,n}$ from $y_{k,n}$ with $p_{\theta_{\mathbf{G}}}(x\mid y_{k,n})$ and evaluate the prompt with ${\mathbf{D}}$ to obtain an in-iteration score $\texttt{Score}^{\prime}(x_{k,n},y_{k,n})=\mathbf{D}(x_{k,n},\hat{x}_{k,n})$. At the end of the iteration, we use the generated $y_{k,n}$ and its score to update $p_{\theta_{\mathbf{F}}}(y\mid x)$. After the entire process, we collect the subset of $\{y_{c}\}_{c=1}^{C}$ generated throughout this process that has the $C$-best in-iteration scores. Then we re-evaluate this subset with the total score $\sum_{i=1}^{M}\texttt{Score}(x_{i},y_{c})$ and return the prompt with the best total score. If there is a tie, then we return the prompt with the highest log likelihood [1]. The pseudocode for the algorithm is outlined in Algorithm 1 and we also illustrate this algorithm in Figure 2.
The key difference between PRISM  and prior methods is that PRISM  updates the entire sampling distribution of prompts, whereas prior works [6, 39, 16] directly update the tokens of a single prompt or the embedding of the prompt. We believe that maintaining the whole prompt distribution is beneficial as text-to-image generation is not a one-to-one operation, i.e. an image can be described by multiple different text prompts and the same text prompt can correspond to multiple different generated images. Having access to the whole distribution allows the method to sample a more diverse range of prompts without starting from scratch and may also help the optimization escape potential local optima.
Since PRISM only requires samples from ${\mathbf{G}}$ conditioned on the prompts, one may use any T2I generative model of their choice. On the other hand, more careful treatment is required for designing ${\mathbf{F}}$ and ${\mathbf{D}}$. We will elaborate on these design decisions below.
@@ 3.3 Designing and updating $\mathbf{F}$ and $p_{\theta_{\mathbf{F}}}$
3.3.1 What is $p(y\mid x)$?
In general, it is not obvious what the joint or the conditional distribution of all text and images is, so some form of approximation is unavoidable. In the context of image generation, a natural choice of the image-conditioned text distribution is an image captioning model. Traditional captioning models, however, fall short in controlled image generation for two primary reasons: (1) The level of detail necessary for generating specific images far exceeds what generic captioning models provide [13]; (2) effective prompts for T2I models are often not grammatically correct sentences but rather collection of phrases that describe the details about the image, which generic captioning models are not trained to generate. For example, in Figure 5, the second reference image is generated by the prompt “A broken robot lays on the ground with plants growing over it, somber, HD, hyper realistic, intricate detail” with Stable Diffusion, but a caption for this image will not include components like “HD” or “hyper realistic”. As a result, instead of “a good description of an image”, we wish to directly model “possible prompts that are used to generate this image”.
3.3.2 Desiderata
@@
3.3.3 Multimodal LLM
stands out as the ideal choice for ${\mathbf{F}}$ due to their ability to directly tailor the generation of prompts via system prompts and to adapt through in-context learning without requiring access to the model’s parameters. Specifically, since the model can ingest both images and texts, we can incorporate the reference images, intermediate prompts and generated images, and the score associated with the generated images all in the context of the LLM. Then, the model can be prompted to jointly reason over all available information and perform in-context learning. The in-context learning facilitates iterative refinement of the prompt to update the posterior distribution based on feedback or even additional human instructions, without the need for model retraining. Concretely, the model would process how the image generative model is affected by different prompts, propose improvements, and create new prompts, much like a prompt engineer. More precisely, in practice, we design system prompts that explicitly condition the LLM to generate improvements and new prompts given the results from the previous iterations, similar to the chain-of-thought [38] technique.

### v2

@@ Abstract
Prompt engineering is an effective but labor-intensive way to control text-to-image (T2I) generative models. Its time-intensive nature and complexity have spurred the development of algorithms for automated prompt generation. However, these methods often struggle with transferability across T2I models, require white-box access to the underlying model, or produce non-intuitive prompts. In this work, we introduce PRISM, an algorithm that automatically produces human-interpretable and transferable prompts that can effectively generate desired concepts given only black-box access to T2I models. Inspired by large language model (LLM) jailbreaking, PRISM leverages the in-context learning ability of LLMs to iteratively refine the candidate prompt distribution built upon the reference images. Our experiments demonstrate the versatility and effectiveness of PRISM in generating accurate prompts for objects, styles, and images across multiple T2I models, including Stable Diffusion, DALL-E, and Midjourney.
@@ 1 Introduction
An important goal of generative modeling is to design algorithms capable of steering generative models to produce desired output images. Early attempts, which often centered on particular architectures or tasks, were largely characterized by manually-curated data collection, fine-tuning, or retraining from scratch (Srivastava & Salakhutdinov, 2012; Mirza & Osindero, 2014; Isola et al., 2017; Zhu et al., 2017). These requirements are often costly, and the resulting solutions usually do not transfer well between models. Thus despite the promise of these methods, efficient and generalized algorithms for controllable generation remain sought after.
@@
Today, perhaps the most popular approach for controllable generation is to guide the generation process with a piece of textual information, or prompt, that describes the properties of the desired output using text-to-image (T2I) generative models (Rombach et al., 2022; Yu et al., 2022). Through text, T2I models allow users to quickly and easily describe a wide variety of concepts, and users can more efficiently explore the behavior of their model through a myriad of strategies Chao et al. (2023); Wen et al. (2023). The predominant method for obtaining such input text is to manually design candidate prompts in an iterative, trial-and-error fashion, a process known as prompt engineering, based on what the user (prompt engineer) believes will lead to a desirable output. Unfortunately, these practices are often sensitive to different phrasings (Webson & Pavlick, 2022), require expert domain knowledge, and are notably inefficient as they necessitate a human in the loop.
@@
Motivated by the drawbacks of manual prompt engineering, a recent line of work known as personalized or subject-driven T2I generation has sought to automate the controllable generation pipeline. Given a collection of reference images that capture specific concepts, such as artistic style or shared objects, personalized T2I algorithms are designed to produce images that reflect those concepts illustrated in the reference images. While personalized T2I methods often involve fine-tuning or retraining the underlying T2I model (Ruiz et al., 2023; Chen et al., 2023; Shi et al., 2023), several approaches focus specifically on automating prompt engineering to generate effective prompts. Unfortunately, existing algorithms in this spirit tend to require pre-collected, architecture-specific keywords [^1] https://github.com/pharmapsychotic/clip-interrogator or white-box, embedding-based optimization (Gal et al., 2023; Mahajan et al., 2023), leading to non-interpretable prompts (Wen et al., 2023) and preclude the possibility of directly generating prompts for closed-source T2I models (e.g., Midjourney or DALL-E).
@@
In order to address these shortcomings, we propose Prompt Refinement and Iterative Sampling Mechanism (PRISM), a new automated prompt engineering algorithm for personalized T2I generation. A key observation is that prompt engineers repeat the process of updating their “belief” of what makes an effective prompt based on the difference between their desired results and the generated images from previous iterations. Inspired by jailbreaking attacks on large language models (LLMs) (Chao et al., 2023) and LLMs as optimizers (Pryzant et al., 2023), we design an algorithm that operates with only limited human input, is capable of generating human interpretable and editable prompts, makes minimal assumptions about the underlying T2I model, and generalizes across different T2I models, including popular black-box models such as DALL-E and Midjourney.
@@
Given a set of reference images, our method first generates an initial prompt and its corresponding image using a vision-language model (VLM) as “prompt engineer assistant” and a T2I generator. We then obtain a score indicating the visual similarity of the generated image and the reference image with respect to the targeting concept via another VLM as judge. Leveraging LLMs’ in-context learning abilities (Shin et al., 2020; Zhou et al., 2023; Yang et al., 2024), we instruct the prompt engineer assistant VLM to update the candidate prompt distribution based on the previously generated prompt, images, and the evaluation scores. This processing, shown in Figure 1, is then repeated for a predetermined number of iterations. In the end, PRISM outputs the best-performing prompt by re-evaluating the top prompts generated from this process. In this way, PRISM seamlessly integrates iterative reasoning into the image generation process, much like a real prompt engineer. Our approach can therefore go beyond basic image-to-image transformations and conventional single-shot methods, providing a more versatile and robust framework for generating images that are both visually precise and contextually relevant.
@@
Experimentally, our method shows significantly better generalizability and transferability as we achieve the best performance in almost all metrics when experimenting with closed-source models in comparison to baselines including Textual Inversion (Gal et al., 2023), PEZ (Wen et al., 2023), BLIP2 (Li et al., 2023) and CLIP-Interrogator1. Our results also indicate that PRISM consistently outperforms existing methods with respect to human-interpretability while maintaining high visual accuracy. Finally, we demonstrate that the strong human interpretability makes the prompts generated by PRISM easily editable, unlocking a wide array of creative possibilities in real life.
Figure 1: Given a set of reference images, our method, PRISM, is capable of creating human-interpretable and accurate prompts for the desired concept that are also transferable to both open-sourced and closed-sourced text-to-image models. $\bigoplus$ denotes prompt concatenation.
@@ 2 Related Works
Controllable T2I generation
Several methods tackle conditional image generation in a training-free manner by using pretrained diffusion models as priors (Meng et al., 2022; Chung et al., 2023; Song et al., 2022; He et al., 2023), and analogous approaches exist for T2I diffusion models (Yu et al., 2023; Rout et al., 2023; He et al., 2024). However, these methods assume that the controllability objectives can be formulated as differentiable loss functions, require access to model parameters and involve complex hyperparameter tuning. Another class of approaches (Zhang et al., 2023; Ye et al., 2023; Ruiz et al., 2023; Chen et al., 2023; Shi et al., 2023) also improve the controllability of pretrained T2I models, but they require expensive fine-tuning or re-training of the underlying model every time they are applied to a new task. Prompt tuning methods (Gal et al., 2023; Wen et al., 2023; Mahajan et al., 2023) are in the same spirit as this paper, as they do not require training of the T2I model and condition generations on given reference images. However, unlike PRISM, these methods require access to the underlying model parameters or produce non-interpretable prompts.
Prompt engineering
Manual prompt engineering is a popular approach to eliciting desired behaviors from large pre-trained models because it uses little or no data and does not require fine-tuning (Radford et al., 2021; Brown et al., 2020). However, major drawbacks of manual prompt engineering include its laborious nature, its reliance on domain expertise, and its sensitivity to phrasings (Lu et al., 2022; Webson & Pavlick, 2022). To address this issue, several methods have been proposed to construct prompts in an automated manner (Shin et al., 2020; Gao et al., 2021; Zhou et al., 2022; 2023; Manikandan et al., 2023; Pryzant et al., 2023; Yang et al., 2024), and some have applied similar techniques to various downstream tasks (Mañas et al., 2024; Liu et al., 2024; Yang et al., 2023; Hao et al., 2024). In particular,  Liu et al. (2024) applied the algorithm they designed for image classification to image inversion. Moreoever, LLM jailbreaking focuses on automatically designing prompts that elicits specific content (often objectionable or illicit) from a targeted LLM (Zou et al., 2023; Wei et al., 2024; Robey et al., 2023; Liu et al., 2023). A particularly relevant work is Chao et al. (2023), which uses an auxiliary LLM to iteratively construct jailbreak prompts. Our method builds on this idea to create prompts to generate images satisfying the desired criteria.
@@ 3 Method
Figure 2: An illustration of PRISM. “System” indicates the system prompt setups for the VLMs.
@@ 3.1 Problem Statement
First, let $x\in{\mathcal{X}}$ denote an image, and $y\in{\mathcal{Y}}$ denote a textual prompt. Given a collection of reference images $\{x_{i}\}_{i=1}^{M}$, a prompt engineer $\mathbf{F}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ samples a candidate prompt $y$ corresponding to each reference image $x$, i.e., $y\sim p_{\theta_{{\mathbf{F}}}}(y\mid x)$. A T2I generative model $\mathbf{G}:{\mathcal{Y}}\rightarrow{\mathcal{X}}$ then uses this candidate prompt to generate a new image, $x\sim p_{\theta_{\mathbf{G}}}(x\mid y)$, and a judge model $\mathbf{D}:{\mathcal{X}}\times{\mathcal{X}}\rightarrow[0,1]$ then scores the visual similarity between the images based on some criteria. Our goal is then to find the best prompt:
$y^{\star}\left(\{x_{i}\}_{i=1}^{M}\right)=\operatorname*{arg\,max}_{y\in{\mathcal{Y}}}\sum_{i=1}^{M}s(x_{i},y),$ (1)
where $s(x_{\text{target}},y)=\mathbb{E}_{x\sim p_{\theta_{\mathbf{G}}}(x\mid y)}\left[{\mathbf{D}}(x,x_{\text{target}})\right]$.
@@ 3.2 Algorithm
8:         $\mathbf{D}$ calculates an in-iteration score $s^{\prime}(x_{k,n},y_{k,n})=\mathbf{D}(x_{k,n},\hat{x}_{k,n})$
9:         Update $p_{\theta_{\mathbf{F}}}$ based on $x_{k,n},\hat{x}_{k,n},y_{k,n},s^{\prime}(x_{k,n},y_{k,n})$ and the chat history of stream $n$
@@
13:Re-evaluate this subset with total score $\sum_{i=1}^{M}s(x_{i},y_{c})$
14:Return the prompt with the best total score. In case of a tie, return the prompt with the highest log likelihood.
Our method, Prompt Refinement and Iterative Sampling Mechanism  (PRISM), is an iterative process that repeats a prompt refinement subroutine for $K$ iterations in $N$ parallel streams, where $N\times K$ is a predetermined compute budget. At iteration $k$, the $n$-th stream of PRISM  randomly selects a reference image $x_{k,n}$ from $\{x_{i}\}_{i=1}^{M}$ and uses ${\mathbf{F}}$ to sample a candidate prompt $y_{k,n}$ from $p_{\theta_{\mathbf{F}}}(y\mid x_{k,n})$. Then it queries ${\mathbf{G}}$ to generate a single $\hat{x}_{k,n}$ from $y_{k,n}$ with $p_{\theta_{\mathbf{G}}}(x\mid y_{k,n})$ and evaluate the prompt with ${\mathbf{D}}$ to obtain an in-iteration score $s^{\prime}(x_{k,n},y_{k,n})=\mathbf{D}(x_{k,n},\hat{x}_{k,n})$. At the end of the iteration, we use the generated $y_{k,n}$ and its score to update $p_{\theta_{\mathbf{F}}}(y\mid x)$. After the entire process, we collect the subset of $\{y_{c}\}_{c=1}^{C}$ generated throughout this process that has the $C$-best in-iteration scores. Then we re-evaluate this subset with the total score $\sum_{i=1}^{M}s(x_{i},y_{c})$ and return the prompt with the best total score. If there is a tie, then we return the prompt with the highest log likelihood. The pseudocode and an illustration for this algorithm are outlined in Algorithm 1 and Figure 2 respectively.
The key difference between PRISM  and prior methods is that PRISM  updates the entire sampling distribution of prompts, whereas prior works (Gal et al., 2023; Wen et al., 2023; Mahajan et al., 2023) directly update the tokens of a single prompt or the embeddings of the prompt. We believe that maintaining the whole prompt distribution is beneficial as text-to-image generation is not a one-to-one operation, i.e. an image can be described by multiple different text prompts and the same text prompt can correspond to multiple differently generated images. Having access to the whole distribution allows the method to sample a more diverse range of prompts without starting from scratch and may also help the optimization escape potential local optima.
Since PRISM only requires samples from ${\mathbf{G}}$, one may use any T2I model of their choice. However, careful consideration is needed when designing ${\mathbf{F}}$ and ${\mathbf{D}}$, which we will elaborate on below.
@@ 3.3 Designing and updating $\mathbf{F}$ and $p_{\theta_{\mathbf{F}}}$
What is $p(y\mid x)$?
In general, it is not obvious what the joint or the conditional distribution of all text and images is, so some form of approximation is unavoidable. In the context of image generation, a natural choice of the image-conditioned text distribution is an image captioning model. Traditional captioning models, however, fall short in controlled image generation for two primary reasons: (1) The level of detail necessary for generating specific images far exceeds what generic captioning models provide (Liang et al., 2023); (2) effective prompts for T2I models are often not grammatically correct sentences but rather collection of phrases that describe the details about the image, which generic captioning models are not trained to generate. For example, in Figure 5, the second reference image is generated by the prompt “A broken robot lays on the ground with plants growing over it, somber, HD, hyper realistic, intricate detail” with Stable Diffusion, but a caption for this image will not include components like “HD” or “hyper realistic”. As a result, instead of “a good description of an image”, we wish to directly model “possible prompts that are used to generate this image”.
Desiderata
@@
Vision-Language Models as ${\mathbf{F}}$
VLMs stand out as the ideal choice for ${\mathbf{F}}$ due to their ability to directly tailor the generation of prompts via system prompts and to adapt through in-context learning without requiring access to the model’s parameters. Specifically, since the model can ingest both images and texts, we can incorporate the history of reference images, intermediate prompts, generated images, and the evaluation scores all in the context of the LLM. Then, the model can be prompted to jointly reason over all available information and perform in-context learning. The in-context learning facilitates iterative refinement of the prompt to update the posterior distribution based on feedback or even additional human instructions, without the need for model retraining. Concretely, the model would process how the image generative model is affected by different prompts, propose improvements, and create new prompts, much like a prompt engineer. This way, we can naturally incorporate iterative reasoning into the image generation process and go beyond simple image-to-image transformations and traditional single-shot generation, and thus offers a more robust and versatile framework for producing accurate and contextually relevant images. In practice, we design system prompts that explicitly condition the LLM to generate improvements and new prompts given the results from the previous iterations, similar to the chain-of-thought (Wei et al., 2022) and textual gradients (Pryzant et al., 2023) technique.